OpenStudy (anonymous):

GIVING MEDAL Find csc theta if the terminal side of theta passes through the point (-4,3).

OpenStudy (solomonzelman):

 right?

OpenStudy (anonymous):

Yes @SolomonZelman

OpenStudy (solomonzelman):

so you can tell already that 

OpenStudy (solomonzelman):

now you need to find the hypotenuse, using \[a^2+b^2=c^2\] plug in your numbers,\[3^3+4^2=c^2\]\[25=c^2\]\[c=5\]

OpenStudy (solomonzelman):

so, you know that the opposite side is 3 and the hypotenuse is 5. Sin = opp / hyp and csc is the inverse of sin, so Csc = hyp / opp --> in your case, it's 5/3.